OpenStudy (anonymous):

Carl can paint a room 3 hours faster than Jennifer can. If they work together, they can complete the job in 2 hours. How long does it take Jennifer to complete this job on her own?

OpenStudy (anonymous):

Why would you multiply the first rate by 3?

OpenStudy (phi):

typo! Here is the corrected version et Jennifer paint 1 room in T hours. Her rate is 1 room/T hours or 1/T rooms per hour Carl does 1 room in T-3 hours, so his rate is 1/(T-3) now use rate * time = 1 room in 2 hours they finish a room together: \[ \frac{1}{T}\cdot 2 + \frac{1}{T-3}\cdot 2 = 1 \] now solve for T to find how long it takes for Jennifer to paint a room